# and knot coloring invariants

### v2

@@ Abstract.
Generalized Legendrian racks are nonassociative algebraic structures based on the Legendrian Reidemeister moves. We study algebraic aspects of GL-racks and coloring invariants of Legendrian links.
We answer an open question characterizing the group of GL-structures on a given rack. As applications, we classify several infinite families of GL-racks. We also compute automorphism groups of dihedral GL-quandles and the categorical center of GL-racks.
Then we construct an equivalence of categories between racks and GL-quandles.
@@
Answering another open question, we use GL-racks to distinguish Legendrian knots whose classical invariants are identical. In particular, we complete the classification of Legendrian $8_{13}$ knots.
@@
Algebraic theory, center, coloring invariant, Legendrian knot, medial quandle, rack, tensor product
@@ 1. Introduction
Generalized Legendrian racks, also called GL-racks or bi-Legendrian racks, are a nonassociative algebraic structure used to distinguish Legendrian links in $\mathbb{R}^{3}$ or $S^{3}$. GL-racks can be traced back to algebraic structures called kei, which Takasaki [takasaki] introduced in 1942 to study symmetric spaces; quandles, which Joyce [joyce] introduced in 1982 to study conjugation in groups and links in $\mathbb{R}^{3}$ and $S^{3}$; and racks, which Fenn and Rourke [fenn] introduced in 1992 to study framed links in $3$-manifolds. Kei, quandles, and racks have enjoyed significant study as link invariants in geometric topology and in their own rights in quantum algebra and group theory.
More recently, various authors have introduced variants of racks suitable for studying Legendrian links. In 2017, Kulkarni and Prathamesh [original] introduced rack invariants of Legendrian knots. In 2021, Ceniceros et al. [ceniceros] refined these invariants by introducing Legendrian racks. In 2023, Karmakar et al. [karmakar] and Kimura [bi] independently strengthened these constructions by introducing GL-racks, which are racks equipped with additional structure.
In this paper, we further develop the theory and applications of GL-racks, answer open questions about GL-racks and Legendrian knot classifications, and classify GL-racks up to order $8$. In particular, we provide the first examples of GL-racks that distinguish Legendrian knots that are not distinguishable by their classical invariants, answering a question of Kimura [bi]*Section 4.
@@ 1.1. The structure of this paper
Sections 46 study GL-racks from group-theoretic, categorical, and universal-algebraic perspectives, respectively; readers only interested in applications to Legendrian knot theory may safely skip these sections. In Section 4, we determine all GL-structures on a given rack, answering a question of Karmakar et al. [karmakar]*Section 3. As applications, we classify all GL-structures on permutation racks, conjugation quandles of abelian and centerless groups, Takasaki kei of abelian $2$-torsion-free groups, and dihedral quandles. We also characterize GL-rack automorphism groups and compute them for dihedral quandles. Finally, we compute the centers of the categories of GL-racks, GL-quandles, Legendrian racks, and Legendrian quandles.
In Section 5, we construct an equivalence of categories between the categories of racks and GL-quandles. This equivalence restricts to an equivalence between the full subcategories of medial racks and medial GL-quandles. Our construction corresponds to an isomorphism of algebraic theories.
In Section 6, we study tensor products of racks and GL-racks coming from universal algebra. We show that the categories of racks and GL-racks have tensor units. This is unusual for noncommutative algebraic theories; for example, the tensor product in the category of groups does not have a tensor unit. We show that the induced symmetric monoidal structure on medial racks is closed, and similarly for medial GL-racks.
In Section 7, we discuss the applications of GL-racks to Legendrian knot theory. We use GL-rack coloring invariants to distinguish several conjecturally distinct Legendrian knots with the same classical invariants. This gives a positive answer to a question of Kimura [bi]*Section 4; we also show in Appendix A.3 that Legendrian racks give a positive answer. In particular, we add to the classification of Legendrian $8_{10}$ knots and complete the classification of Legendrian $8_{13}$ knots in the extended Legendrian knot atlases of Bhattacharyya et al. [atlas-site] and Petkova and Schwartz [new-atlas].
In Section 8, we propose questions for further research based on our results.
In Appendix A, we describe algorithms that can classify finite GL-racks of a given order up to isomorphism. These algorithms use Vojtěchovský and Yang’s [library] classification of racks up to order $11$. We provide implementations of our algorithms in GAP [GAP4] and the data we were able to compute and enumerate for all $n\leq 8$. We also provide an algorithm that computes GL-rack coloring invariants of Legendrian links. As an application, we show that Legendrian racks positively answer the question of Kimura [bi]*Section 4 answered in Section 7 for GL-racks.
@@ Acknowledgments.
I am indebted to Sam Raskin, my undergraduate thesis adviser, for his mentorship, support, and many insights throughout the research and writing processes. I thank Samantha Pezzimenti and Wing Hong Tony Wong for introducing me to Legendrian knot theory and mathematics research, Jose Ceniceros for informing me of the categorical literature on racks, and Peyton Wood for introducing me to GL-racks. I also thank Noam Scully for helpful discussions, Qiaochu Yuan for a terminological correction, and an anonymous Stack Exchange user for referencing me to [alg-theories]*Theorem 3.10.3.
This paper was written in partial fulfillment of the requirements for the bachelor’s degree in mathematics at Yale. I thank Patrick Devlin, Miki Havlíčková, Kati Hubley, Matthew King, Andrew Neitzke, and Carol Rutschman for inspiring me to pursue and complete this degree.
@@ 2.1.2. Racks and quandles
Example 2.2.
[quandlebook]*Example 99 Let $X$ be a set, and fix $\sigma\in S_{X}$. Define $s:X\rightarrow S_{X}$ by $x\mapsto\sigma$, so that $s_{x}(y)=\sigma(y)$ for all $x,y\in X$. Then $(X,\sigma)_{\mathrm{perm}}:=(X,s)$ is a rack called a permutation rack or constant action rack. (Our notation $(X,\sigma)_{\mathrm{perm}}$ is nonstandard, unlike the notation in the next three examples.) Note that $(X,\sigma)_{\mathrm{perm}}$ is a quandle if and only if $\sigma=\operatorname{id}_{X}$, in which case we call $(X,\operatorname{id}_{X})_{\mathrm{perm}}$ a trivial quandle.
Example 2.3.
[quandlebook]*Example 54 Let $A$ be an abelian additive group. Define $s:A\to S_{A}$ by $b\mapsto s_{b}$ with $s_{b}(a):=2b-a$ for all elements $a,b\in A$. Then $T(A):=(A,s)$ is a quandle called a Takasaki kei. Takasaki kei are the earliest examples of racks in the literature; Takasaki [takasaki] introduced them in 1943 to study symmetric spaces.
@@
Example 2.5.
[book]*Example 2.13 Let $X$ be a union of conjugacy classes in a group $G$, and define $c^{G}:X\rightarrow S_{X}$ by sending $x$ to the conjugation map $c^{G}_{x}:=[y\mapsto xyx^{-1}]$. Then $\operatorname{Conj}X:=(X,c^{G})$ is a quandle called a conjugation quandle or conjugacy quandle.
@@
Example 2.7.
[quandlebook]*Example 66 Let $n\geq 3$ be an integer. We call the Takasaki kei $R_{n}:=T(\mathbb{Z}/n\mathbb{Z})$ a dihedral quandle due to the following observation: if $\Sigma=\{s,rs,\dots,r^{n-1}s\}$ is the set of all reflections in the dihedral group $D_{n}=\langle r,s\mid r^{n}=s^{2}=1,\,srs=r^{-1}\rangle$ of order $2n$, then the map $\varphi:\mathbb{Z}/n\mathbb{Z}\to\Sigma$ defined by $k\mapsto r^{k}s$ is a rack isomorphism from $R_{n}$ to $\operatorname{Conj}\Sigma$.
@@ 2.1.3. Dual racks
Definition 2.8.
[dual]*Definition 2.6 Given a rack $R=(X,s)$, define $s^{\prime}:X\to S_{X}$ by $x\mapsto s_{x}^{-1}$. Then $R^{\operatorname{op}}:=(X,s^{\prime})$ is a rack called the dual rack of $R$. (Our notation $R^{\operatorname{op}}$ is nonstandard.)
@@ 2.1.4. Medial racks
Thus, $\mathsf{Rack}$ is complete and cocomplete (see [alg-theories]*Theorem 3.4.5), as we will need later. In particular, the Cartesian product of racks has a natural rack structure, so we can define the following.
Definition 2.11.
[rack-roll]*Section 3 A rack $(X,s)$ is called medial or abelian if the map $X\times X\to X$ defined by $(x,y)\mapsto s_{y}(x)$ is a rack homomorphism. Equivalently, the subgroup $\langle s_{x}s_{y}^{-1}\mid x,y\in X\rangle$ of $\operatorname{Inn}X$ is abelian; see, for example, [medial-quandles]*Proposition 2.1. This subgroup is called the transvection group or displacement group of $(X,s)$. Equivalently, for all $x,y,z\in X$,
@@ 2.2.1. Definition of $\theta_{R}$
Given a rack $R=(X,s)$, define $\theta_{R}:X\to X$ by ${x\mapsto s_{x}(x)}$; see [center]*Proposition 2.5.
@@ 2.2.3. Properties of $\theta$
Hence, $\Theta$ generates a cyclic subgroup of $Z(\mathsf{Rack})$. In fact, this subgroup is $Z(\mathsf{Rack})$; see [center]*Theorem 5.4.
@@ 2.3.1. Examples of quotient racks
Recall that $\mathsf{Rack}$ is the category of models of the algebraic theory of racks in $\mathsf{Set}$, so $\mathsf{Rack}$ is complete and cocomplete. Thus, we can consider quotients of racks by congruences; see [alg-theories]*Lemma 3.5.1 and cf. [dual]*Proposition 3.6. This allows us to state the following definitions, which are actually left adjoints of the inclusion functors from $\mathsf{Qnd}$ and $\mathsf{Rack}_{\mathrm{med}}$ into $\mathsf{Rack}$.
Definition 2.18.
[fenn]*Section 2, [joyce]*Section 10 Given a rack $R=(X,s)$, the associated quandle of $R$ is the quandle $R_{\mathrm{qnd}}:=R/{\sim}$, where $\sim$ is the smallest congruence on $X$ such that $s_{x}(x)=x$ for all $x\in X$. Similarly, the medialization or abelianization of $R$ is the medial rack $R_{\mathrm{med}}:=R/{\sim}$, where $\sim$ is the smallest congruence on $X$ such that
@@ 2.3.2. Free racks
In analogy to free groups or free modules, the free rack $\operatorname{Free}_{\mathsf{Rack}}(X)$ on a set $X$ exists and is uniquely (up to isomorphism) characterized by a universal property; see, for example, [alg-theories]*Corollary 3.7.8. Namely, there exists a set map $i:X\to\operatorname{Free}_{\mathsf{Rack}}(X)$ such that for all racks $R=(Y,t)$ and all set maps $\varphi_{0}:X\to Y$, there exists a unique rack homomorphism ${\varphi\in\operatorname{Hom}_{\mathsf{Rack}}(\operatorname{Free}_{\mathsf{Rack}}(X),R)}$ such that $\varphi i=\varphi_{0}$. That is, the following diagram commutes:
@@
For an explicit set-theoretic construction of free racks, see [free-z]*Proposition 1.3. We can also define the free quandle on $X$ to be $(\operatorname{Free}_{\mathsf{Rack}}(X))_{\mathrm{qnd}}$; cf. [fenn]*Section 7.1.
Example 2.19.
[free-z]*Example 1.6 The free rack $F$ on one generator is canonically isomorphic to the permutation rack $(\mathbb{Z},\sigma)_{\mathrm{perm}}$, where $\sigma(k)=k+1$ for all $k\in\mathbb{Z}$. Under this identification, $\theta_{F}=\sigma$. By contrast, the free quandle on one generator is the trivial quandle with one element.
@@ 2.3.3. Commutative algebraic theories
Medial racks are the largest commutative subtheory of racks. Recall that an algebraic theory $\mathcal{T}$ is called commutative if, for all $\mathcal{T}$-models $A$, every $n$-ary operation $\alpha:\mathcal{T}^{n}\to\mathcal{T}$ defines a homomorphism $A^{n}\to A$; see [alg-theories]*Theorem 3.10.3.
Let $\mathcal{T}_{\operatorname{ab}}$ be the largest commutative subtheory of an algebraic theory $\mathcal{T}$. If $\mathcal{C}$ is the category of $\mathcal{T}$-models in $\mathsf{Set}$, let $\mathcal{C}_{\operatorname{ab}}$ be the full subcategory of $\mathcal{C}$ whose objects are $\mathcal{T}_{\operatorname{ab}}$-models. The following strengthens part of [alg-theories]*Theorem 3.10.3; we could not find a reference for this precise statement.
@@
We refer the reader to the proof of the statement in [alg-theories]*Theorem 3.10.3 that (1) implies (3). This proof uses the commutativity of $n$-ary operations in $Y$ but not in $X$, so the implication still holds with the weakened assumption that $X$ is not necessarily a $\mathcal{T}_{\operatorname{ab}}$-model. ∎
@@ 2.3.4. Mediality and Hom racks
Hence, Proposition 2.20 recovers the following result of Grøsfjeld [rack-roll]*Proposition 3.3 in 2021, which in turn generalized a result of Crans and Nelson [Hom]*Theorem 3 in 2014.

### v3

# Classification and structure of generalized Legendrian racks

@@ Abstract.
We study algebraic aspects of generalized Legendrian racks, which are nonassociative structures based on the Legendrian Reidemeister moves.
We answer an open question characterizing the group of GL-structures on a given rack. As applications, we classify several infinite families of GL-racks. We also compute automorphism groups of dihedral GL-quandles.
Then we compute the centers of the category of GL-racks and several of its full subcategories. We also construct an equivalence of categories between racks and GL-quandles.
@@
Algebraic theory, center, equivalence of categories, Legendrian knot invariant, medial, quandle, rack, tensor product
@@ 1. Introduction
Generalized Legendrian racks, also called GL-racks or bi-Legendrian racks, are a nonassociative algebraic structure used to distinguish Legendrian links in $\mathbb{R}^{3}$ or $S^{3}$. GL-racks can be traced back to algebraic structures called kei, which Takasaki [takasaki] introduced in 1942 to study symmetric spaces; quandles, which Joyce [joyce] and Matveev [matveev] independently introduced in 1982 to study links in $\mathbb{R}^{3}$ and $S^{3}$; and racks, which Fenn and Rourke [fenn] introduced in 1992 to study framed links in $3$-manifolds. Kei, quandles, and racks have enjoyed significant study as link invariants in geometric topology and in their own rights in quantum algebra and group theory.
More recently, various authors have introduced variants of racks suitable for studying Legendrian links. In 2017, Kulkarni and Prathamesh [original] introduced rack invariants of Legendrian knots. In 2021, Ceniceros et al. [ceniceros] refined these invariants by introducing Legendrian racks. In 2023, Karmakar et al. [karmakar] and Kimura [bi] independently strengthened these constructions by introducing GL-racks, which are racks equipped with additional structure. After completing the initial preprint of this paper, the author [ta] also studied involutory GL-racks in relation to similar algebraic structures used to study surface-knots.
In this paper, we further develop the theory and classification of GL-racks. In particular, we study categorical and group-theoretic aspects of GL-racks, and we classify GL-racks up to order $8$.
@@ 1.1. The structure of this paper
In Section 4, we study group-theoretic aspects of GL-racks. We determine all GL-structures on a given rack, answering a question of Karmakar et al. [karmakar]*Sec. 3. As applications, we classify all GL-structures on permutation racks, certain conjugation quandles, and certain Takasaki kei, including all dihedral quandles. We also characterize GL-rack automorphism groups and compute them for dihedral quandles equipped with GL-structures.
In Section 5, we study categorical aspects of GL-racks. First, we compute the centers of the categories of GL-racks, GL-quandles, Legendrian racks, and Legendrian quandles. Then we construct an equivalence of categories between the categories of racks and GL-quandles. This equivalence restricts to an equivalence between the full subcategories of medial racks and medial GL-quandles. Our construction corresponds to an isomorphism of algebraic theories.
In Section 6, we study tensor products of racks and GL-racks coming from universal algebra. We show that the categories of racks and GL-racks have tensor units. This is unusual for noncommutative algebraic theories; for example, the tensor product in the category of groups does not have a tensor unit. The induced symmetric monoidal structure on medial racks is closed, and similarly for medial GL-racks.
In Section 7, we propose questions for further research.
In Appendix A, we describe algorithms that can classify finite GL-racks of a given order up to isomorphism. These algorithms use Vojtěchovský and Yang’s [library] classification of racks up to order $11$. We provide implementations of our algorithms in GAP [GAP4] and the data we were able to compute and enumerate for all $n\leq 8$.
@@ Acknowledgments.
I thank Sam Raskin for providing mentorship, support, and insights throughout the research and writing processes. I thank Samantha Pezzimenti, Tony W. H. Wong, Jose Ceniceros, and Peyton Wood for introducing me to various aspects of the theory, and I thank Noam Scully for helpful discussions. I also thank Zhiyi He for identifying errors in an earlier draft; these errors were entirely my own. The affected section has been omitted. I thank Qiaochu Yuan for a terminological correction and an anonymous Stack Exchange user for referencing me to [alg-theories].
This research was conducted in partial fulfillment of the senior requirements for the bachelor’s degree in mathematics at Yale University. I thank Patrick Devlin, Miki Havlíčková, Kati Hubley, Matthew King, Andrew Neitzke, and Carol Rutschman for inspiring me to complete this degree.
@@ 2.1.2. Racks and quandles
Example 2.2 ([quandlebook]*Ex. 99).
Let $X$ be a set, and fix $\sigma\in S_{X}$. Define $s:X\rightarrow S_{X}$ by $x\mapsto\sigma$, so that $s_{x}(y)=\sigma(y)$ for all $x,y\in X$. Then $(X,\sigma)_{\mathrm{perm}}:=(X,s)$ is a rack called a permutation rack or constant action rack. (Our notation $(X,\sigma)_{\mathrm{perm}}$ is nonstandard, unlike the notation in the next three examples.) Note that $(X,\sigma)_{\mathrm{perm}}$ is a quandle if and only if $\sigma=\operatorname{id}_{X}$, in which case we call $(X,\operatorname{id}_{X})_{\mathrm{perm}}$ a trivial quandle.
Example 2.3 ([quandlebook]*Ex. 54).
Let $A$ be an abelian additive group. Define $s:A\to S_{A}$ by $b\mapsto s_{b}$ with $s_{b}(a):=2b-a$ for all elements $a,b\in A$. Then $T(A):=(A,s)$ is a quandle called a Takasaki kei. Takasaki kei are the earliest examples of racks in the literature; Takasaki [takasaki] introduced them in 1943 to study symmetric spaces.
@@
Example 2.5 ([book]*Ex. 2.13).
Let $X$ be a union of conjugacy classes in a group $G$, and define $c^{G}:X\rightarrow S_{X}$ by sending $x$ to the conjugation map $c^{G}_{x}:=[y\mapsto xyx^{-1}]$. Then $\operatorname{Conj}X:=(X,c^{G})$ is a quandle called a conjugation quandle or conjugacy quandle.
@@
Example 2.7 ([quandlebook]*Ex. 66).
Let $n\geq 3$ be an integer. We call the Takasaki kei $R_{n}:=T(\mathbb{Z}/n\mathbb{Z})$ a dihedral quandle due to the following observation: if $\Sigma=\{s,rs,\dots,r^{n-1}s\}$ is the set of all reflections in the dihedral group $D_{n}=\langle r,s\mid r^{n}=s^{2}=1,\,srs=r^{-1}\rangle$ of order $2n$, then the map $\varphi:\mathbb{Z}/n\mathbb{Z}\to\Sigma$ defined by $k\mapsto r^{k}s$ is a rack isomorphism from $R_{n}$ to $\operatorname{Conj}\Sigma$.
@@ 2.1.3. Dual racks
Definition 2.8 ([dual]*Def. 2.6).
Given a rack $R=(X,s)$, define $s^{\prime}:X\to S_{X}$ by $x\mapsto s_{x}^{-1}$. Then $R^{\operatorname{op}}:=(X,s^{\prime})$ is a rack called the dual rack of $R$. (Our notation $R^{\operatorname{op}}$ is nonstandard.)
@@ 2.1.4. Medial racks
Thus, $\mathsf{Rack}$ is complete and cocomplete (see [alg-theories]*Thm. 3.4.5), as we will need later. In particular, the Cartesian product of racks has a natural rack structure, so we can define the following.
Definition 2.11 ([rack-roll]*Sec. 3).
A rack $(X,s)$ is called medial or abelian if the map $X\times X\to X$ defined by $(x,y)\mapsto s_{y}(x)$ is a rack homomorphism. Equivalently, the subgroup $\langle s_{x}s_{y}^{-1}\mid x,y\in X\rangle$ of $\operatorname{Inn}X$ is abelian; see, for example, [medial-quandles]*Prop. 2.1. This subgroup is called the transvection group or displacement group of $(X,s)$. Equivalently, for all $x,y,z\in X$,
@@ 2.2.1. Definition of $\theta_{R}$
Given a rack $R=(X,s)$, define $\theta_{R}:X\to X$ by ${x\mapsto s_{x}(x)}$; see [center]*Prop. 2.5.
@@ 2.2.3. Properties of $\theta$
Hence, $\Theta$ generates a cyclic subgroup of $Z(\mathsf{Rack})$. In fact, this subgroup is $Z(\mathsf{Rack})$; see [center]*Thm. 5.4.
@@ 2.3.1. Examples of quotient racks
Recall that $\mathsf{Rack}$ is the category of models of the algebraic theory of racks in $\mathsf{Set}$, so $\mathsf{Rack}$ is complete and cocomplete. Thus, we can consider quotients of racks by congruences; see [alg-theories]*Lem. 3.5.1 and cf. [dual]*Prop. 3.6. This allows us to state the following definitions, which are actually left adjoints of the inclusion functors from $\mathsf{Qnd}$ and $\mathsf{Rack}_{\mathrm{med}}$ into $\mathsf{Rack}$.
Definition 2.18 ([fenn]*Sec. 2, [joyce]*Sec. 10).
Given a rack $R=(X,s)$, the associated quandle of $R$ is the quandle $R_{\mathrm{qnd}}:=R/{\sim}$, where $\sim$ is the smallest congruence on $X$ such that $s_{x}(x)=x$ for all $x\in X$. Similarly, the medialization or abelianization of $R$ is the medial rack $R_{\mathrm{med}}:=R/{\sim}$, where $\sim$ is the smallest congruence on $X$ such that
@@ 2.3.2. Free racks
In analogy to free groups or free modules, the free rack $\operatorname{Free}_{\mathsf{Rack}}(X)$ on a set $X$ exists and is uniquely (up to isomorphism) characterized by a universal property; see, for example, [alg-theories]*Cor. 3.7.8. Namely, there exists a set map $i:X\to\operatorname{Free}_{\mathsf{Rack}}(X)$ such that for all racks $R=(Y,t)$ and all set maps $\varphi_{0}:X\to Y$, there exists a unique rack homomorphism ${\varphi\in\operatorname{Hom}_{\mathsf{Rack}}(\operatorname{Free}_{\mathsf{Rack}}(X),R)}$ such that $\varphi i=\varphi_{0}$. That is, the following diagram commutes:
@@
For an explicit set-theoretic construction of free racks, see [free-z]*Prop. 1.3. We can also define the free quandle on $X$ to be $(\operatorname{Free}_{\mathsf{Rack}}(X))_{\mathrm{qnd}}$; cf. [fenn]*Sec. 7.1.
Example 2.19 ([free-z]*Ex. 1.6).
The free rack $F$ on one generator is canonically isomorphic to the permutation rack $(\mathbb{Z},\sigma)_{\mathrm{perm}}$, where $\sigma(k)=k+1$ for all $k\in\mathbb{Z}$. Under this identification, $\theta_{F}=\sigma$. By contrast, the free quandle on one generator is the trivial quandle with one element.
@@ 2.3.3. Commutative algebraic theories
Medial racks are the largest commutative subtheory of racks. Recall that an algebraic theory $\mathcal{T}$ is called commutative if, for all $\mathcal{T}$-models $A$, every $n$-ary operation $\alpha:\mathcal{T}^{n}\to\mathcal{T}$ defines a homomorphism $A^{n}\to A$; see [alg-theories]*Thm. 3.10.3.
Let $\mathcal{T}_{\operatorname{ab}}$ be the largest commutative subtheory of an algebraic theory $\mathcal{T}$. If $\mathcal{C}$ is the category of $\mathcal{T}$-models in $\mathsf{Set}$, let $\mathcal{C}_{\operatorname{ab}}$ be the full subcategory of $\mathcal{C}$ whose objects are $\mathcal{T}_{\operatorname{ab}}$-models. The following strengthens part of [alg-theories]*Thm. 3.10.3; we could not find a reference for this precise statement.
@@
We refer the reader to the proof of the statement in [alg-theories]*Thm. 3.10.3 that (1) implies (3). This proof uses the commutativity of $n$-ary operations in $Y$ but not in $X$, so the implication still holds with the weakened assumption that $X$ is not necessarily a $\mathcal{T}_{\operatorname{ab}}$-model. ∎
@@ 2.3.4. Mediality and Hom racks
Hence, Proposition 2.20 recovers the following result of Grøsfjeld [rack-roll]*Prop. 3.3 in 2021, which in turn generalized a result of Crans and Nelson [Hom]*Thm. 3 in 2014.